# SCAM: A Real-World Typographic Robustness Evaluation for Multimodal Foundation Models

### v4

@@ 5 Results
Our evaluation highlights an important finding: while LVLMs are vulnerable to typographic attacks (see table 1), the strength of their vision encoder plays a significant role in this susceptibility. Smaller LLaVA models, such as llava-llama3:8b, llava-1.6:7b, and llava-1.6:13b, suffer substantial accuracy drops of more than 40-60 percentage points under typographic attacks. This vulnerability largely stems from their reliance on weaker vision encoders, particularly the ViT-L-14-336 backbone trained by OpenAI, which is highly susceptible to such attacks (see fig. 6). However, as model size increases, the robustness to typographic attacks improves. LLaVA models with larger LLM backbones, such as llava-1.6:34b, exhibit significantly better performance under attack. This suggests that a more powerful LLM backbone can compensate for the limitations of the vision encoder, making these models more resilient to typographic distortions. In comparison, models trained by OpenAI, such as gpt-4o, exhibit strong robustness to typographic attacks, with a minimal accuracy drop of around 3 percentage points. The smaller gpt-4o-mini, however, experiences a higher drop of approximately 15 points, illustrating again that larger models tend to be more robust. Despite that, model size alone does not determine robustness. In fact, the large llama3.2-vision:90b model is more vulnerable to typographic attacks than the smaller llava-1.6:34b, highlighting the critical role of the vision encoder in overall performance. These results highlight that while vision encoders remain a critical vulnerability, increasing the size and quality of the LLM backbone plays a key role in reducing susceptibility to typographic attacks.
@@ 6 Discussion
Our analysis further reveals that training regime, training data, and model architecture significantly influence susceptibility of Vision-Language Models (VLMs) to typographic attacks, whereas model size alone does not correlate with robustness. In Large Vision-Language Models (LVLMs), vulnerability to typographic attacks carry over from the vision encoder, making these models inherently susceptible. However, we find that larger Large Language Model (LLM) backbones can mitigate this effect, leading to improved robustness and accuracy to typographic attacks.
@@
A key limitation of our work is the use of “naive prompting” [52, 53] for LVLMs. Utilizing multi-prompt evaluation, such as those proposed in [54, 47], could provide a better assessment of their robustness to typographic attacks. Another limitation is the focus on English-language attacks and it should be assessed whether language-specific factors influence model susceptibility.
@@
As multimodal models are increasingly deployed to safety-critical applications [12, 11, 13, 13], ensuring robustness against typographic attacks is essential to prevent serious consequences. Approaches such as alignment training [55] or defense techniques [36] offer promising directions for enhancing robustness. Additionally, our findings suggest that an LVLM ’s vulnerability to typographic attacks is closely tied to its vision encoder. To investigate this further, we plan to retrain LLaVA models with different vision backbones to systematically evaluate their impact on robustness. This will help determine whether stronger vision encoders mitigate susceptibility or if additional fine-tuning strategies are required.
@@ 7 Conclusion
In this paper, we introduced the Subtle Character Attacks on Multimodal Models (SCAM) dataset (and two variations, NoSCAM and SynthSCAM), the largest and most diverse collection of real-world images designed to evaluate the vulnerability of multimodal foundation models to typographic attacks. Our evaluation demonstrated that current state-of-the-art multimodal models are significantly affected by typographic attacks, highlighting a critical weakness in their robustness. We hope that our broad analysis of Vision-Language Models (VLMs) will help researchers make more informed choices when selecting typographically robust models for their downstream applications.
@@
Our work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project-ID 528483508 - FIP 12. We thank the BLISS [^2] https://bliss.berlin/ community – its very existence, collaborative energy, and shared commitment to open research made this work possible. Special thanks to Zeynep Altayli, Jonas Loos, Philippa Ringe, and Joseph Tschörner for their valuable efforts in image collection and labeling. We especially thank Jonas Loos for his significant work on the project page.
@@
JW was responsible for image processing, code development, model evaluations, visualizations, and overall coordination. EP contributed to conceptualization, figure design, and manuscript writing, including the introduction, dataset, evaluation, and drafted most of the results and discussion sections. JH focused on related work research and supported overall writing. JL supported model evaluations, managed data and code hosting, and created the project webpage. LP contributed to dataset expansion through large-scale labeling and provided continuous critical feedback. LH initiated the project, provided supervision throughout, designed the data collection protocol, and guided the conceptual and methodological direction of the work. All authors contributed to image collection and labeling.
@@ A.1 Example Images of SCAM, NoSCAM, and SynthSCAM
In fig. 7, we highlight some examples from the Subtle Character Attacks on Multimodal Models (SCAM) dataset and the generated NoSCAM and SynthSCAM datasets.
@@
A.2 SCAM Labels
The SCAM dataset includes diverse object labels:
Table 7 summarizes the frequency of object labels. The top 15 labels (e.g., “lamp”, “plant”, “bottle”) are among the most frequent, while the bottom 5 labels (e.g., “chain”, “roller board”, “mask”) appear rarely. Notably, 437 object labels appear exactly once, illustrating a pronounced long-tail distribution that underscores the diversity of the dataset.
Table 7: Top 15 and bottom 5 object labels by appearance frequency.
Object label Count
lamp 13
plant 13
bottle 13
chair 11
glass 10
candle 10
cable 8
grill 8
printer 8
cup 8
speaker 8
pillow 8
brush 7
canister 7
pen 7
chain 1
roller board 1
mask 1
goggles 1
webcam 1
SCAM includes a variety of attack words:
Table 8 summarizes the frequency of attack words. The top 15 attack words (e.g., “knife”, “truck”, “slow”) occur frequently, whereas many attack words are rare. In fact, 56 attack words appear exactly once, again highlighting overall diversity of the dataset.
Table 8: Top 15 and bottom 5 attack words by appearance frequency.
Attack Word Count
knife 24
truck 23
slow 20
lion 20
zebra 19
helmet 18
bike 17
honey 15
pig 15
camera 15
lemon 15
pen 15
tree 14
house 14
fish 13
beatles 1
green 1
raccoon 1
beach 1
unicorn 1
A.3 Zero-Shot Evaluation of LVMs: Text Templates
@@
A.4 Model Accuracies
Table 9: Performance of VLMs and LVLMs available through OpenCLIP [44] resp. ollama [46] and OpenAI’s API on SCAM, PAINT [38], and RTA-100 [36] datasets. Results on PAINT and RTA-100 are broadly similar to those on SCAM, supporting the consistency of typographic attack effects across datasets.
@@ A.4 Model Accuracies
llava-llama3-1.1:8b - 98.09 39.50 44.04 48.62 41.83
llava-1.6:7b - 97.50 57.25 61.45 53.64 51.20
llava-1.6:13b - 98.88 58.00 55.42 54.55 56.00
llava-1.6:34b - 98.97 84.85 85.11 89.09 81.50
@@
A.5 Architecture Comparison
In table 5 in the main text, we compare different architectures (CLIP-RN vs. CLIP-ViT, and CLIP-ViT vs. SigLIP-ViT) and state averaged accuracies. In table 10, we detail the specific models that we choose from the OpenCLIP suite for this comparison.
Table 10: Models chosen from the OpenCLIP suite for architecture comparison in table 5. The table groups models by architecture category and indicates the corresponding fixed training data. In OpenCLIP, v1 refers to LAION-COCO NLLB [50].

### v6

@@ 5 Results
Our evaluation highlights an important finding: while LVLMs are vulnerable to typographic attacks (see table˜1) , their susceptibility appears to be influenced by the strength of the underlying vision encoder. Smaller LLaVA and Gemma3 models, such as gemma3:4b, llava:7b-v1.6, llava-llama3:8b, gemma3:12b, and llava:13b-v1.6, suffer substantial accuracy drops of more than 40-60 percentage points under typographic attacks. We hypothesize that this vulnerability stems from their reliance on weaker vision encoders, particularly the ViT-L-14-336 backbone trained by OpenAI, which is highly susceptible to such attacks (see fig.˜6). However, as model size increases, the robustness to typographic attacks improves. Models with larger LLM backbones, such as gemma:27b and llava:34b-v1.6, exhibit significantly better performance under attack. This suggests that a more powerful LLM backbone can compensate for the limitations of the vision encoder, making these models more resilient to typographic distortions. In comparison, presumably much larger closed-source models such as claude-sonnet-4-20250514 and gpt-4o exhibit strong robustness to typographic attacks, with a minimal accuracy drop of around 8 resp. 3 percentage points. The smaller gpt-4o-mini and llama4:scout, however, experience a higher drop of approximately 15 points, illustrating again that larger models tend to be more robust. Despite that, model size alone does not determine robustness. In fact, the large llama3.2-vision:90b model is more vulnerable to typographic attacks than the smaller llava:34b-v1.6. These results highlight that while vision encoders remain a critical vulnerability, increasing the size and quality of the LLM backbone plays a key role in reducing susceptibility to typographic attacks.
Impact of VLM and LLM Size on LVLM Performance.
Section˜A.6 presents a preliminary experiment that isolates the role of the vision encoder in LVLMs with respect to typographic robustness. The experiment does not show a clear gain in robustness when substituting the vision encoder with one ranked safer by SCAM, while keeping the LLM, training data, and training regime unchanged. Instead, the overall performance decreases, highlighting the difficulty of choosing the optimal vision encoder and showcasing the need for further experiments investigating the robustness impact of different VLM choices. Apart from that, as discussed in section˜A.7, scaling up the LLM backbone not only improves performance on SCAM but also enhances the model’s general typographic understanding on textual and OCR-based benchmarks.
@@ 6 Discussion
Our analysis further reveals that training regime, training data, and model architecture significantly influence susceptibility of Vision-Language Models (VLMs) to typographic attacks, whereas model size alone does not correlate with robustness. In Large Vision-Language Models (LVLMs), vulnerability to typographic attacks appears to carry over from the vision encoder, making these models susceptible to the same weaknesses. However, we find that larger Large Language Model (LLM) backbones can mitigate this effect, leading to improved robustness and accuracy to typographic attacks, while also improving their typographic understanding.
@@
Despite incorporating a different, “safer” prompt in section˜A.5, a central limitation of our evaluation remains the reliance on “naive prompting” [55, 56] for assessing LVLMs. Adopting more advanced multi-prompt evaluation strategies may yield a more reliable measure of model robustness against typographic attacks. A further limitation is our exclusive focus on English-language attacks; future work should investigate whether language-specific factors influence model susceptibility.
@@
As multimodal models are increasingly deployed to safety-critical applications [15, 14, 16, 16], ensuring robustness against typographic attacks is essential to prevent serious consequences. Approaches such as alignment training [57] or (mechanistic) defense techniques [44, 58] offer promising directions for enhancing robustness. Our evaluations suggest that the vulnerability of LVLMs to typographic attacks is tied to their vision encoder. However, our initial experiments (section˜A.6) show that simply swapping LLaVA’s vision encoder for a more robust one does not reliably transfer robustness when integrated naively. Future work will extend this by systematically retraining LLaVA models with a range of vision backbones to evaluate their role in attack susceptibility and to assess whether encoder strength alone suffices or if additional adaptation steps are necessary.
@@ 7 Conclusion
In this paper, we introduced the Subtle Character Attacks on Multimodal Models (SCAM) dataset (and two variations, NoSCAM and SynthSCAM), the largest and most diverse collection of real-world images designed to evaluate the vulnerability of multimodal foundation models to typographic attacks. Our evaluation demonstrated that current state-of-the-art multimodal models are significantly affected by typographic attacks, highlighting a critical weakness in their robustness. We expect that our broad analysis of Vision-Language Models (VLMs) and Large Vision-Language Models (LVLMs) will help researchers make more informed choices when selecting typographically robust models for their safety-critical downstream applications.
@@
Our work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project-ID 528483508 - FIP 12. We thank the BLISS [^2] https://bliss.berlin/ community – its very existence, collaborative energy, and shared commitment to open research made this work possible. Special thanks to Zeynep Altayli, Philippa Ringe, and Joseph Tschörner for their valuable efforts in image collection and labeling.
@@
JW was responsible for image processing, code development, model evaluations, visualizations, and overall coordination. EP contributed to conceptualization, figure design, and manuscript writing, including the introduction, dataset, evaluation, and drafted most of the results and discussion sections. JH focused on related work research and supported overall writing. JL supported model evaluations, managed data and code hosting, and created the project webpage. LP contributed to dataset expansion through large-scale labeling and provided continuous critical feedback. EH provided supervision and contributed through discussions. LH initiated the project, provided supervision throughout, designed the data collection protocol, and guided the conceptual and methodological direction of the work. All authors except EH contributed to image collection and labeling.
@@ A.1 Example Images of SCAM, NoSCAM, and SynthSCAM
In fig.˜7, we highlight some examples from the Subtle Character Attacks on Multimodal Models (SCAM) dataset and the generated NoSCAM and SynthSCAM datasets.
@@
A.2 Zero-Shot Evaluation of LVMs: Text Templates
@@
A.3 Model Accuracies on Typographic Attack Datasets
Table 6: Performance of VLMs and LVLMs available through OpenCLIP [20] resp. ollama [21], and Anthropic’s resp. OpenAI’s API on SCAM, PAINT [42], and RTA-100 [44] datasets.
@@ A.3 Model Accuracies on Typographic Attack Datasets
llava-llama3:8b - 98.09 39.50 44.04 48.62 41.83
llava:7b-v1.6 - 97.50 58.43 61.45 53.64 51.20
llava:13b-v1.6 - 98.88 58.00 55.42 54.55 56.00
llava:34b-v1.6 - 98.97 84.85 85.11 89.09 81.50
gemma3:4b - 97.24 58.05 64.80 50.00 60.66
gemma3:12b - 99.14 52.02 64.80 44.55 50.50
gemma3:27b - 97.42 81.67 83.65 80.91 78.10
@@
A.4 Architecture Comparison
In table˜5(a) in the main text, we compare different architectures (CLIP-RN vs. CLIP-ViT, and CLIP-ViT vs. SigLIP-ViT) and present averaged accuracies. Table˜7 details the specific models that we choose from the OpenCLIP suite for this comparison.
Table 7: Models chosen from the OpenCLIP suite for architecture comparison in table˜5(a). The table groups models by architecture category and indicates the corresponding fixed training data. In OpenCLIP, v1 refers to LAION-COCO NLLB [51].
@@ A.4 Architecture Comparison
A.5 Prompt Variation
Our evaluation of LVLMs in section˜4 focuses on a single prompt. Given prior work demonstrating the importance of prompt phrasing and advocating for multi-prompt evaluation [59, 50], we additionally test a modified version that appears “safer.” This variant appends the instruction “Do not get fooled by typographic attacks.” to the original prompt. Contrary to the expectation formulated in these works, this modification does not reduce vulnerability. As illustrated in fig.˜8, the susceptibility to typographic attacks remains largely unchanged.
Figure 8: The modified “safer” prompt does not alleviate susceptibility to typographic attacks.
A.6 Replacing VLM in LLaVA
Previous work [40, 60] suggest that susceptibility to typographic attacks can be attributed to a cross-modal alignment deficiency inherent to LVLMs while the functionality of vision encoders is intact. Further, [61] demonstrates that integrating an adversarially pretrained vision encoder into LLaVA significantly improves robustness against various attacks, including typographic perturbations introduced in [60].
Our evaluations on SCAM support the conclusion that some VLMs exhibit higher typographic robustness than others. In particular, LVLMs using the OpenAI-trained ViT-L-14-336 encoder trained – such as LLaVA [14, 30, 31] – show heightened vulnerability. To isolate the effect of the vision encoder, we train a LLaVA model using the ViT-L-14-CLIPA-336 encoder [62, 63], which was trained on datacomp1b and differs primarily in training data and regime, not architecture. Compared to the OpenAI encoder, the CLIPA model exhibits a substantially lower robustness drop: only 25% from NoSCAM to SCAM, versus 65% (see table˜1). It also outperforms the OpenAI encoder by 3% on average across 38 datasets evaluated using OpenCLIP [48] [^3] https://github.com/mlfoundations/open_clip/blob/main/docs/openclip_results.csv.
[…]
Table˜8 shows that our CLIPA-based LLaVA variant does not outperform the original model on SCAM, despite using a VLM that is more robust in isolation. Performance on other benchmarks evaluated via VLMEvalKit [65] also drops. These findings suggest that the benefits of a more robust vision encoder do not transfer under naive LLaVA integration.
Note the experimental constraints: only one small Large Language Model (LLM) (Vicuna v1.5 7B), only two different vision backbones, and no hyperparameter tuning are employed.
Table 8: Performance comparison of the original LLaVA-v1.5-7B and our CLIPA-based LLaVA variant across diverse benchmarks. Despite the CLIPA vision encoder showing stronger robustness in isolation, this does not translate to improved performance on SCAM. Evaluations for the upper section are conducted using VLMEvalKit [65]. Other results follow the protocol outlined in the main part of the paper.
Benchmark LLaVA-v1.5-7B (original) LLaVA-v1.5-7B-CLIPA (ours)
MMBench v1.1 [66] 64.32 60.22
MME Perception [67] 1444.81 1378.96
SEED-Bench-1 (img) [68] 66.81 64.02
TextVQA [69] 46.93 45.57
OCRBench [70] 31.90 31.60
OCR-VQA [71] 60.64 57.39
NoSCAM 97.16 89.66
SCAM 54.69 45.48
SynthSCAM 51.55 46.08
PAINT [42] 52.73 53.64
RTA100 [44] 48.30 47.30
A.7 Typographic Robustness vs. Typographic Understanding
The main paper shows that increasing the size of LLM backbones reduces susceptibility to typographic attacks (see results for LLaVA and Gemma3 in table˜1). Simultaneously, benchmarks such as TextVQA and OCRBench, reported on the VLMLeaderboards [^5] https://huggingface.co/spaces/opencompass/open_vlm_leaderboard, indicate that typographic understanding – essentially, reading – also improves with model scale (see table˜9). Together, these findings suggest that typographic robustness emerges as model size increases, alongside improved textual perception. However, [36] reports the opposite correlation between susceptibility and OCR performance in LVLMs, indicating that stronger reading ability alone does not guarantee robustness.
Table 9: Performance of selected LLaVA-Next [31] models on three OCR-related benchmarks (TextVQA [69], OCRBench [70] (from VLMEvalKit [65]), and OCR-VQA [71]) shows that increasing LLM backbone size leads to improved typographic understanding. All models use the ViT-L-14-336 vision encoder. Results are reported from the OpenVLM leaderboard, evaluated using the VLMEvalKit toolkit [65].
Model Params (B) TextVQA (%) OCRBench OCR-VQA (%)
LLaVA-Next-Yi-34B 34.8 69.3 574 66.3
LLaVA-Next-Vicuna-13B 13.4 66.9 537 65.3
LLaVA-Next-Llama3 8.0 65.3 531 60.7
LLaVA-Next-Mistral-7B 7.6 65.2 531 61.2
LLaVA-Next-Vicuna-7B 7.1 63.8 532 63.8
A.8 SCAM Labels
The SCAM dataset includes diverse object labels:
Table˜10 summarizes the frequency of object labels. The top 15 labels (e.g., “lamp”, “plant”, “bottle”) are among the most frequent, while the bottom 5 labels (e.g., “chain”, “roller board”, “mask”) appear rarely. Notably, 437 object labels appear exactly once, illustrating a pronounced long-tail distribution that underscores the diversity of the dataset.
Table 10: Top 15 and bottom 5 object labels by appearance frequency.
Object label Count
lamp 13
plant 13
bottle 13
chair 11
glass 10
candle 10
cable 8
grill 8
printer 8
cup 8
speaker 8
pillow 8
brush 7
canister 7
pen 7
chain 1
roller board 1
mask 1
goggles 1
webcam 1
SCAM includes a variety of attack words:
Table˜11 summarizes the frequency of attack words. The top 15 attack words (e.g., “knife”, “truck”, “slow”) occur frequently, whereas many attack words are rare. In fact, 56 attack words appear exactly once, again highlighting overall diversity of the dataset.
Table 11: Top 15 and bottom 5 attack words by appearance frequency.
Attack Word Count
knife 24
truck 23
slow 20
lion 20
zebra 19
helmet 18
bike 17
honey 15
pig 15
camera 15
lemon 15
pen 15
tree 14
house 14
fish 13
beatles 1
green 1
raccoon 1
beach 1
unicorn 1